An astronaut travels at a speed of 0.98c from the
Answer #1

1 light year = 9.46x10^15

6.0x10^17 m = 6.0x10^17 / 9.46x10^15 = 63.4249 light years

63.4249 light years away

t = 63.4249/0.98 = 64.72

T = t * sqrt(1-v^2/c^2)

T = 64.72 x sqrt(1-0.98^2/1^2)

T = 12.87 years

34 +12.87 = 46.9years old

She is 46.9 or 47 years old when she arrives at the star.